# Mastering AI: Big Data, Deep Learning, and the Evolution of Large Language Models - Blockchain and Applications

### v2

@@ 19.3.2 In-Game Items and Virtual Real Estate
A popular example is the game Axie Infinity[222], where players collect and battle digital creatures (called Axies) that are represented as NFTs. These creatures can be bought, sold, or bred, with some rare Axies fetching high prices.
Another example is Decentraland[339], a virtual world where users can purchase plots of virtual land as NFTs, which they can develop, sell, or lease to other users.
@@ 19.3.3 Digital Copyright for Music and Multimedia
For example, the musician 3LAU[340] made headlines by selling a series of NFTs representing digital music albums, generating millions of dollars in sales. This model allows creators to bypass traditional music platforms and directly engage with fans, while also receiving royalties from subsequent sales.
@@ 20.1 Basic Concepts of DAO
Decentralized Autonomous Organizations (DAOs) are one of the most groundbreaking innovations enabled by blockchain technology[219]. They represent a new form of organization that operates without centralized control, relying on smart contracts and a decentralized governance structure. In this chapter, we will explore the fundamental concepts of DAOs, including their governance structure, voting mechanisms, and practical applications in both governance and decentralized finance (DeFi).
@@ 20.1.1 Decentralized Governance Structure of DAOs
At the heart of a DAO is its decentralized governance structure. Unlike traditional organizations, where decisions are typically made by a central authority or a board of directors, DAOs distribute decision-making power among the members[341]. This is achieved through the ownership of tokens, which act as voting shares in the organization.
The core of this decentralized governance is often implemented through smart contracts[342]. A smart contract is a self-executing contract where the terms of the agreement are written directly into code. In a DAO, smart contracts define the rules of the organization and automatically enforce them without the need for intermediaries.
For example, a DAO may have a rule that requires a majority vote for any decision to be implemented. Token holders can propose changes or new initiatives, and voting is conducted transparently on the blockchain[343]. Once a proposal receives the required number of votes, the smart contract automatically enforces the decision, whether it’s allocating funds, changing a governance parameter, or implementing new features.
@@ 20.1.2 Voting Mechanisms in DAOs
In many DAOs, voting power is proportional to the number of tokens held by a member. This is known as token-weighted voting[344]. The more tokens you own, the greater your influence on the outcome of a decision. While this system is straightforward, it can lead to centralization, where large token holders have disproportionate control.
@@
Quadratic voting is a more complex system that aims to balance the influence of participants[345]. Instead of voting power being directly proportional to the number of tokens held, quadratic voting allows members to allocate votes in a way that reflects the intensity of their preferences. The cost of casting additional votes increases quadratically, making it expensive for large token holders to dominate the process.
@@ 20.2.1 DAO Applications in Governance Projects
DAOs are commonly used in governance projects to allow communities to manage decentralized protocols, allocate funds, and make collective decisions. A well-known example of this is the MakerDAO[220], which governs the Maker Protocol, a decentralized lending platform that manages the stablecoin DAI[346].
@@
  • Uniswap DAO: Governs the Uniswap decentralized exchange protocol[237]. The community votes on proposals to adjust fees, add new features, or upgrade the protocol.
  • Compound DAO: Manages the Compound lending protocol[347], allowing the community to propose and vote on protocol changes.
  • Gitcoin DAO: Focuses on funding open-source projects through decentralized governance and community-driven funding decisions[348].
@@ 20.2.2 The Role of DAOs in DeFi
DAOs play a crucial role in the decentralized finance (DeFi) ecosystem. Many DeFi protocols are governed by DAOs, which manage everything from protocol upgrades to the allocation of treasury funds[349]. By allowing token holders to vote on key decisions, DAOs enhance the transparency and decentralization of these platforms.
For instance, in the Aave protocol[217], a DAO governs the allocation of its funds, decisions regarding new token listings, and the parameters of the lending pools. This decentralized governance ensures that no single entity controls the platform, and decisions are made by the community in a transparent manner.
@@ 20.3.1 Legal Status and Regulatory Challenges of DAOs
One of the biggest challenges facing DAOs is their legal status[350]. Since DAOs operate without a centralized entity, it is often unclear who is liable for decisions made by the organization. In traditional organizations, legal entities like corporations or LLCs provide a structure for accountability and liability. However, in DAOs, the decentralized nature makes it difficult to assign responsibility.
Some DAOs have attempted to address this by incorporating as legal entities, such as Wyoming’s DAO LLC law[351], which provides a legal framework for DAOs to operate as LLCs. However, this is still a developing area, and there are many uncertainties regarding how DAOs will be treated by regulators in different jurisdictions.
@@ 20.3.2 Future Directions of DAO Development
One of the key areas of development is improving governance models[219, 352]. While token-weighted voting is the most common system today, new models like quadratic voting and reputation-based voting are being explored to create fairer and more inclusive decision-making processes.
Scalability is another challenge that DAOs need to address. As DAOs grow, the complexity of decision-making increases, and it becomes harder to reach consensus[353]. Solutions such as layer-2 scaling solutions[189] and cross-chain governance mechanisms[354] are being developed to address these issues.
@@ 21.1.1 Distributed Ledger Technology
Distributed Ledger Technology (DLT)[142] is the core innovation behind blockchain, providing a decentralized method for storing data across multiple nodes in a network. The key features of DLT include security, decentralization, and scalability, which are crucial for the success of blockchain systems.
@@
As blockchain systems grow, one of the main challenges is ensuring scalability—the ability to handle an increasing number of transactions without degrading performance. For example, in Bitcoin’s blockchain, every node must validate each transaction, which limits throughput to around 7 transactions per second (TPS)[174]. Solutions like sharding[232] divide the ledger into smaller partitions, allowing parallel processing of transactions.
@@
Security in DLT involves protecting the ledger from malicious actors who might attempt to alter transaction history. The immutability of blockchain, where data once written cannot be changed, provides strong protection[355]. However, researchers continue to explore enhanced cryptographic techniques like quantum-resistant algorithms[356], which aim to safeguard against future threats posed by quantum computing.
@@
Decentralization ensures that no single entity controls the ledger. This is a defining feature of blockchain and prevents issues like censorship or single points of failure. However, true decentralization introduces the scalability trilemma[357]—the challenge of balancing decentralization, security, and scalability simultaneously, which continues to be a major research focus.
@@
A notable example of a blockchain system with improved DLT architecture is Ethereum 2.0[230], which uses a combination of Proof of Stake (PoS) and sharding to increase scalability while maintaining decentralization.
@@ 21.1.2 Optimization and Improvement of Consensus Algorithms
The consensus algorithm in a blockchain is the method by which the network reaches agreement on the validity of transactions. Research in this area focuses on optimizing existing algorithms like Proof of Work (PoW) and Proof of Stake (PoS), as well as developing new ones[149].
@@
PoW[149], used in Bitcoin, requires miners to solve computational puzzles to validate transactions. This algorithm is secure but highly energy-intensive. Research on ASIC resistance[358] and alternative approaches like Proof of Space[359] aims to mitigate the environmental impact.
@@
PoS[231], used in Ethereum 2.0, selects validators based on the amount of cryptocurrency they hold and are willing to ”stake” as collateral. PoS is more energy-efficient than PoW but introduces new challenges such as ensuring fairness and preventing centralization of wealth.
@@
Emerging consensus mechanisms such as Delegated Proof of Stake (DPoS)[231] and Byzantine Fault Tolerance (BFT) aim to improve scalability and security. For example, DPoS allows token holders to vote for delegates who validate transactions, which increases efficiency but requires trust in elected delegates.
@@
The Tendermint consensus algorithm[360], which is based on BFT, is used by the Cosmos Network[361]. It ensures fast finality, meaning transactions are confirmed almost instantly, addressing the scalability issue present in traditional PoW systems.
@@ 21.1.3 Privacy Protection and Anonymity Research
ZKPs[288] allow a user to prove the validity of a transaction without revealing the transaction details. Zcash [362]is a prominent example of a blockchain that uses zk-SNARKs (Zero-Knowledge Succinct Non-Interactive Arguments of Knowledge[363]) to ensure privacy.
@@
Researchers are exploring methods to execute smart contracts without exposing sensitive data. One approach is using MPC (Multi-Party Computation[364]), where multiple parties jointly compute a function while keeping their inputs private.
@@
While privacy is crucial, blockchain systems must still ensure transparency for regulatory compliance. Researchers are investigating solutions like selective disclosure, where only authorized entities can view certain transaction details[365].
@@
In the Monero[366] blockchain, privacy is enhanced through ring signatures and stealth addresses, which obscure the origin and destination of transactions, making it difficult to trace.
@@ 21.2.1 Integration of Blockchain and Cryptography
Blockchain relies on cryptographic hashing to secure transactions. A hash function takes input data and produces a fixed-size string, which serves as a digital fingerprint of the data. For example, Bitcoin uses the SHA-256 hash function to secure its blocks[367].
@@
Digital signatures authenticate the identity of users. In blockchain, each transaction is signed by the sender using their private key. The recipient can verify this signature using the sender’s public key. The Elliptic Curve Digital Signature Algorithm (ECDSA) is widely used for this purpose in Bitcoin[368].
@@
Data encryption ensures that only authorized parties can view sensitive information. Research on homomorphic encryption[369] allows computations to be performed on encrypted data, which is beneficial for privacy-preserving applications in blockchain.
@@ 21.2.2 Blockchain and Game Theory
Game theory provides a framework to analyze how participants behave in a decentralized system, such as a blockchain. Researchers use game theory to study incentive structures and strategies that ensure participants act in a way that benefits the network[370].
@@
In blockchain systems like Bitcoin, miners are incentivized to validate transactions through rewards. However, game theory helps explore how to prevent malicious behaviors, like selfish mining[371], where miners might collude to increase their rewards.
@@
Game theory is also used to analyze how participants can cooperate to maintain security[372]. For instance, in Proof of Stake systems, validators are penalized for dishonest behavior, which deters attacks.
@@ 21.2.3 Blockchain and Economics
DeFi platforms allow users to borrow, lend, and trade assets without intermediaries like banks. Researchers study how these systems can operate efficiently while ensuring security and preventing fraud[292, 291].
@@
The design of cryptocurrencies and tokens involves economic principles like scarcity and incentive alignment. Researchers explore how tokenomics can create sustainable ecosystems that incentivize participation and growth[373].
@@ 21.3.1 Major Academic Conferences in the Blockchain Field
This annual conference focuses on blockchain technology, covering topics like consensus algorithms, privacy, and applications in various industries[374].
@@
As one of the premier conferences on cryptography, Crypto often includes sessions on blockchain-related cryptographic research[375].
@@
ICBC is another major event where researchers discuss advancements in blockchain infrastructure, scalability, and real-world applications[376].
@@
This journal focuses specifically on blockchain technology, covering a wide range of topics from cryptographic foundations to industry applications.
@@
Ledger[377] is an open-access journal that publishes peer-reviewed research on blockchain and cryptocurrency, with a focus on both technical and economic aspects.
@@
Journals like Journal of Cryptology[378] and Cryptography and Communications[379] also publish papers related to the cryptographic techniques used in blockchain systems.

### v3

@@ 19.3.2 In-Game Items and Virtual Real Estate
A popular example is the game Axie Infinity[axieinfinityAxieInfinity], where players collect and battle digital creatures (called Axies) that are represented as NFTs. These creatures can be bought, sold, or bred, with some rare Axies fetching high prices.
Another example is Decentraland[decenrtalandDecentraland], a virtual world where users can purchase plots of virtual land as NFTs, which they can develop, sell, or lease to other users.
@@ 19.3.3 Digital Copyright for Music and Multimedia
For example, the musician 3LAU[3lau3LAUOfficial] made headlines by selling a series of NFTs representing digital music albums, generating millions of dollars in sales. This model allows creators to bypass traditional music platforms and directly engage with fans, while also receiving royalties from subsequent sales.
@@ 20.1 Basic Concepts of DAO
Decentralized Autonomous Organizations (DAOs) are one of the most groundbreaking innovations enabled by blockchain technology[Ding2023ASO]. They represent a new form of organization that operates without centralized control, relying on smart contracts and a decentralized governance structure. In this chapter, we will explore the fundamental concepts of DAOs, including their governance structure, voting mechanisms, and practical applications in both governance and decentralized finance (DeFi).
@@ 20.1.1 Decentralized Governance Structure of DAOs
At the heart of a DAO is its decentralized governance structure. Unlike traditional organizations, where decisions are typically made by a central authority or a board of directors, DAOs distribute decision-making power among the members[Schneider2022DecentralizedAO]. This is achieved through the ownership of tokens, which act as voting shares in the organization.
The core of this decentralized governance is often implemented through smart contracts[Balcerzak2022BlockchainTA]. A smart contract is a self-executing contract where the terms of the agreement are written directly into code. In a DAO, smart contracts define the rules of the organization and automatically enforce them without the need for intermediaries.
For example, a DAO may have a rule that requires a majority vote for any decision to be implemented. Token holders can propose changes or new initiatives, and voting is conducted transparently on the blockchain[Han2023DAOG]. Once a proposal receives the required number of votes, the smart contract automatically enforces the decision, whether it’s allocating funds, changing a governance parameter, or implementing new features.
@@ 20.1.2 Voting Mechanisms in DAOs
In many DAOs, voting power is proportional to the number of tokens held by a member. This is known as token-weighted voting[Kurniawan2022VotingMS]. The more tokens you own, the greater your influence on the outcome of a decision. While this system is straightforward, it can lead to centralization, where large token holders have disproportionate control.
@@
Quadratic voting is a more complex system that aims to balance the influence of participants[Posner2014VotingSQ]. Instead of voting power being directly proportional to the number of tokens held, quadratic voting allows members to allocate votes in a way that reflects the intensity of their preferences. The cost of casting additional votes increases quadratically, making it expensive for large token holders to dominate the process.
@@ 20.2.1 DAO Applications in Governance Projects
DAOs are commonly used in governance projects to allow communities to manage decentralized protocols, allocate funds, and make collective decisions. A well-known example of this is the MakerDAO[makerdaoMakerDAOUnbiased], which governs the Maker Protocol, a decentralized lending platform that manages the stablecoin DAI[Kozhan2021DecentralizedSA].
@@
  • Uniswap DAO: Governs the Uniswap decentralized exchange protocol[uniswapUniswapInterface]. The community votes on proposals to adjust fees, add new features, or upgrade the protocol.
  • Compound DAO: Manages the Compound lending protocol[compoundCompound], allowing the community to propose and vote on protocol changes.
  • Gitcoin DAO: Focuses on funding open-source projects through decentralized governance and community-driven funding decisions[gitcoinGitcoinGovernance].
@@ 20.2.2 The Role of DAOs in DeFi
DAOs play a crucial role in the decentralized finance (DeFi) ecosystem. Many DeFi protocols are governed by DAOs, which manage everything from protocol upgrades to the allocation of treasury funds[Ushida2021RegulatoryCO]. By allowing token holders to vote on key decisions, DAOs enhance the transparency and decentralization of these platforms.
For instance, in the Aave protocol[aaveAave], a DAO governs the allocation of its funds, decisions regarding new token listings, and the parameters of the lending pools. This decentralized governance ensures that no single entity controls the platform, and decisions are made by the community in a transparent manner.
@@ 20.3.1 Legal Status and Regulatory Challenges of DAOs
One of the biggest challenges facing DAOs is their legal status[Guillaume2023DecentralizedAO]. Since DAOs operate without a centralized entity, it is often unclear who is liable for decisions made by the organization. In traditional organizations, legal entities like corporations or LLCs provide a structure for accountability and liability. However, in DAOs, the decentralized nature makes it difficult to assign responsibility.
Some DAOs have attempted to address this by incorporating as legal entities, such as Wyoming’s DAO LLC law[wyolegLegislativeService], which provides a legal framework for DAOs to operate as LLCs. However, this is still a developing area, and there are many uncertainties regarding how DAOs will be treated by regulators in different jurisdictions.
@@ 20.3.2 Future Directions of DAO Development
One of the key areas of development is improving governance models[Ding2023ASO, Gilson2024EnhancingTD]. While token-weighted voting is the most common system today, new models like quadratic voting and reputation-based voting are being explored to create fairer and more inclusive decision-making processes.
Scalability is another challenge that DAOs need to address. As DAOs grow, the complexity of decision-making increases, and it becomes harder to reach consensus[Sheikh2023BuiltTL]. Solutions such as layer-2 scaling solutions[Mandal2023InvestigatingLS] and cross-chain governance mechanisms[DiRose2018ComparisonAA] are being developed to address these issues.
@@ 21.1.1 Distributed Ledger Technology
Distributed Ledger Technology (DLT)[Gutlapalli2016CommercialAO] is the core innovation behind blockchain, providing a decentralized method for storing data across multiple nodes in a network. The key features of DLT include security, decentralization, and scalability, which are crucial for the success of blockchain systems.
@@
As blockchain systems grow, one of the main challenges is ensuring scalability—the ability to handle an increasing number of transactions without degrading performance. For example, in Bitcoin’s blockchain, every node must validate each transaction, which limits throughput to around 7 transactions per second (TPS)[Gbel2017IncreasedBS]. Solutions like sharding[CortesGoicoechea2020ResourceAO] divide the ledger into smaller partitions, allowing parallel processing of transactions.
@@
Security in DLT involves protecting the ledger from malicious actors who might attempt to alter transaction history. The immutability of blockchain, where data once written cannot be changed, provides strong protection[N2023TheII]. However, researchers continue to explore enhanced cryptographic techniques like quantum-resistant algorithms[Nisha2024LatticeBasedCA], which aim to safeguard against future threats posed by quantum computing.
@@
Decentralization ensures that no single entity controls the ledger. This is a defining feature of blockchain and prevents issues like censorship or single points of failure. However, true decentralization introduces the scalability trilemma[Nakai2023TheBT]—the challenge of balancing decentralization, security, and scalability simultaneously, which continues to be a major research focus.
@@
A notable example of a blockchain system with improved DLT architecture is Ethereum 2.0[Asif2023ShapingTF], which uses a combination of Proof of Stake (PoS) and sharding to increase scalability while maintaining decentralization.
@@ 21.1.2 Optimization and Improvement of Consensus Algorithms
The consensus algorithm in a blockchain is the method by which the network reaches agreement on the validity of transactions. Research in this area focuses on optimizing existing algorithms like Proof of Work (PoW) and Proof of Stake (PoS), as well as developing new ones[ferdous2020blockchainconsensusalgorithmssurvey].
@@
PoW[ferdous2020blockchainconsensusalgorithmssurvey], used in Bitcoin, requires miners to solve computational puzzles to validate transactions. This algorithm is secure but highly energy-intensive. Research on ASIC resistance[Cho2018ASICResistanceOM] and alternative approaches like Proof of Space[Dziembowski2015ProofsOS] aims to mitigate the environmental impact.
@@
PoS[Wendl2022TheEI], used in Ethereum 2.0, selects validators based on the amount of cryptocurrency they hold and are willing to ”stake” as collateral. PoS is more energy-efficient than PoW but introduces new challenges such as ensuring fairness and preventing centralization of wealth.
@@
Emerging consensus mechanisms such as Delegated Proof of Stake (DPoS)[Wendl2022TheEI] and Byzantine Fault Tolerance (BFT) aim to improve scalability and security. For example, DPoS allows token holders to vote for delegates who validate transactions, which increases efficiency but requires trust in elected delegates.
@@
The Tendermint consensus algorithm[Lei2020ChainedTA], which is based on BFT, is used by the Cosmos Network[cosmosCosmosInternet]. It ensures fast finality, meaning transactions are confirmed almost instantly, addressing the scalability issue present in traditional PoW systems.
@@ 21.1.3 Privacy Protection and Anonymity Research
ZKPs[ZHOU2024103678] allow a user to prove the validity of a transaction without revealing the transaction details. Zcash [zZcashPrivacyprotecting]is a prominent example of a blockchain that uses zk-SNARKs (Zero-Knowledge Succinct Non-Interactive Arguments of Knowledge[Nitulescu2020zkSNARKsAG]) to ensure privacy.
@@
Researchers are exploring methods to execute smart contracts without exposing sensitive data. One approach is using MPC (Multi-Party Computation[Pei2018SmartCB]), where multiple parties jointly compute a function while keeping their inputs private.
@@
While privacy is crucial, blockchain systems must still ensure transparency for regulatory compliance. Researchers are investigating solutions like selective disclosure, where only authorized entities can view certain transaction details[Mukta2023ABI].
@@
In the Monero[getmoneroMoneroProject] blockchain, privacy is enhanced through ring signatures and stealth addresses, which obscure the origin and destination of transactions, making it difficult to trace.
@@ 21.2.1 Integration of Blockchain and Cryptography
Blockchain relies on cryptographic hashing to secure transactions. A hash function takes input data and produces a fixed-size string, which serves as a digital fingerprint of the data. For example, Bitcoin uses the SHA-256 hash function to secure its blocks[Courtois2014OptimizingSI].
@@
Digital signatures authenticate the identity of users. In blockchain, each transaction is signed by the sender using their private key. The recipient can verify this signature using the sender’s public key. The Elliptic Curve Digital Signature Algorithm (ECDSA) is widely used for this purpose in Bitcoin[Wang2014SecureIO].
@@
Data encryption ensures that only authorized parties can view sensitive information. Research on homomorphic encryption[Zhang2016ARO] allows computations to be performed on encrypted data, which is beneficial for privacy-preserving applications in blockchain.
@@ 21.2.2 Blockchain and Game Theory
Game theory provides a framework to analyze how participants behave in a decentralized system, such as a blockchain. Researchers use game theory to study incentive structures and strategies that ensure participants act in a way that benefits the network[Liu2019ASO].
@@
In blockchain systems like Bitcoin, miners are incentivized to validate transactions through rewards. However, game theory helps explore how to prevent malicious behaviors, like selfish mining[Li2020RationalPA], where miners might collude to increase their rewards.
@@
Game theory is also used to analyze how participants can cooperate to maintain security[Bhudia2023GameTM]. For instance, in Proof of Stake systems, validators are penalized for dishonest behavior, which deters attacks.
@@ 21.2.3 Blockchain and Economics
DeFi platforms allow users to borrow, lend, and trade assets without intermediaries like banks. Researchers study how these systems can operate efficiently while ensuring security and preventing fraud[Salami2021ChallengesAA, Gupta2022DecentralizedFA].
@@
The design of cryptocurrencies and tokens involves economic principles like scarcity and incentive alignment. Researchers explore how tokenomics can create sustainable ecosystems that incentivize participation and growth[Lamberty2020LeadingDS].
@@ 21.3.1 Major Academic Conferences in the Blockchain Field
This annual conference focuses on blockchain technology, covering topics like consensus algorithms, privacy, and applications in various industries[IEEECon].
@@
As one of the premier conferences on cryptography, Crypto often includes sessions on blockchain-related cryptographic research[iacrCrypto2024].
@@
ICBC is another major event where researchers discuss advancements in blockchain infrastructure, scalability, and real-world applications[ieeeicbcIEEEInternational].
@@
This journal focuses specifically on blockchain technology, covering a wide range of topics from cryptographic foundations to industry applications[].
@@
Ledger[ledgerjournalAboutJournal] is an open-access journal that publishes peer-reviewed research on blockchain and cryptocurrency, with a focus on both technical and economic aspects.
@@
Journals like Journal of Cryptology[springerJournalCryptology] and Cryptography and Communications[springerCryptographyCommunications] also publish papers related to the cryptographic techniques used in blockchain systems.